Felix Company has decided to use a predetermined rate to assign factory overhead to production. The...

  1. Felix Company has decided to use a predetermined rate to assign factory overhead to production. The following predictions have been made for 2019:

Estimate total factory overhead costs

$480,000

Estimate direct labor hours

40,000

DLH

Estimate direct labor costs

$400,000

Estimate machine hours

120,000

MH

Actual direct labor hours

38,000

DLH

Actual direct labor costs

$416,000

Actual machine hours

122,000

MH

Required:

  1. Compute the predetermined factory overhead rate under three different bases: (1) direct labor hours, (2) direct labor costs, and (3) machine hours. (Round amounts to 2 decimal places.)
  1. Assume that actual factory overhead was $506,000 and that Felix Company elected to apply factory overhead to Work in Process based on direct-labor-cost.
  1. Was factory overhead or 2019 over- or under-applied? How much?

11. Give the journal entry to record the application of overhead.

  1. Felix Company follows the policy of writing off any under- or over-applied Factory Overhead balance to Cost of Goods Sold at the end of the year. Give the journal entry necessary at the end of 2019 to dispose of the Factory Overhead balance determined in Part (b).
Answer #1
Felix Company
Part-a(3) : Base on Machine Hour
Predetermined Overhead Rate = (Budgeted factory overhead / Budgeted machine hours)
=480000/120000= $4 per Machine Hour
Part-a(2) : Base on Direct Labour Cost
Predetermined Overhead Rate = (Budgeted factory overhead / Budgeted Direct Labour Cost )
=480000/400000= 120%
Part-a(1) : Base on Direct Labour Hour
Predetermined Overhead Rate = (Budgeted factory overhead / Budgeted Direct Labour Hour )
=480000/40000= 12 Per Labour Hour
Part-b
Applied Overhead based on Direct labour Cost   = Predetermined Overhead Rate X Actual Direct Labour Cost
=120%X 416000= $499200
Part-i
Over Applied /(Under Applied Overhead)= Applied Overhead- Actual Overhead
$499200- $506000= $6800 Underapplied
Journal Entry
S. No. Account Name Debit Credit
1 Factory department overhead control $6,800
Cost of Goods Sold $6,800
(To record underapplied Overhead to cost of good sold)